OpenStudy (anonymous):

Evaluate f^-2g for f = 3 and g = 27

OpenStudy (anonymous):

3^-54

OpenStudy (anonymous):

= 1.72 * 10^-26

OpenStudy (anonymous):

a very tiny number

OpenStudy (anonymous):

i think it is\[f ^{-2}* g\] so it becomes\[g/f ^{2} = 27/3^{2} = 27/9 = 3\]

OpenStudy (anonymous):

have i misread it?

OpenStudy (anonymous):

maybe u r right and i am wrong... only kdrift can tell which interpretation is correct.....